Consider an individual who is five years away from retirement and will need to withdraw all her retirement funds at that time. She has $2,000 in pretax income to allocate each year to a retirement plan, faces a fixed tax rate of 15 percent now as well as at retirement, and anticipates a stable 8 percent return on her investments. She can set up a Roth IRA for a one-time, up-front fee of $10, or she can set up a traditional IRA for free. Which option should she choose?
Answer #1

Solution:

Traditional IRA (Pre-taxed): 2000*(1.08)^6 + 2000*(1.08)^5 + 2000*(1.08) 4 + 2000*(1.08)^3 + 2000*(1.08)^2 + 2000*(1.08)^1 = 15,845.60672

Since the tax rate is 15%, thus will have 15,845.60672 * 85% = 13,468.7657

 

Roth IRA (Post-taxed): 1700*(1.08)^6 + 1700*(1.08)^5 + 1700*(1.08)^4 + 1700*(1.08)^3 + 1700*(1.08)^2 + 1700*(1.08)^1 = 13,468.7657

 

Although the retirement plans will have the same amount (calculated above) of funds available to withdraw when the person retires, however the Roth IRA has a set up fee. Thus must choose the Traditional IRA for avoiding incurring a $10 set up charge